# What Breaks Knowledge Graph based RAG? Empirical Insights into Reasoning under Incomplete Knowledge

### v3

@@ A.1 AMIE3 Candidate Rule Refinement
AMIE3 generates candidate rules by a refinement process using a classical breadth-first search lajus2020fast. It begins with rules that contain only a head atom (e.g. $\Rightarrow\texttt{hasSibling}(X,Y)$) and refines them by adding atoms to the body. For example, it may generate the refined rule:
@@ Appendix B Properties of Horn Rules Mined by AMIE3
These restrictions are widely adopted in KG rule mining (galarraga2015fast; lajus2020fast) to guarantee that discovered rules are logically well-formed and meaningful for downstream reasoning tasks.
@@ B.1 PCA Confidence
Definition. To address this, AMIE3 introduces PCA confidence (Partial Completeness Assumption confidence) (galarraga2015fast), an optimistic variant that partially compensates for KG incompleteness. Given a rule of the form
@@
For further details, see (galarraga2015fast).
@@ B.2 Rule Mining Procedure
AMIE3 generates candidate rules by a refinement process using a classical breadth-first search lajus2020fast. Algorithm 2 summarizes the rule mining process of AMIE3. The algorithm starts with an initial set of rules that contain only a head atom (i.e. $\top\Rightarrow r(X,Y)$, where $\top$ denotes an empty body) and maintains a queue of rule candidates (Line 1). At each step, AMIE3 dequeues a rule $R$ from the queue and evaluates whether it satisfies three criteria (Line 5):
@@ C.1 Recall and Precision
C.2 Evaluation of Generate-on-Graph (GoG) (xu2024generate)
This section provides a detailed analysis of GoG (xu2024generate), following the experimental results presented in the main text. While GoG explicitly addresses knowledge incompleteness as part of its motivation, we did not include this aspect in the main body for several reasons.
GoG’s implementation is heavily optimized for specific datasets, such as CWQ (Talmor2018WSQ) and WebQSP (Yih2016webqsp), and relies on several domain-specific heuristics:
@@ Appendix D Prompt Template
To ensure reproducibility and mitigate randomness in LLM outputs potyka2024robust, we set the generation temperature to 0 in all experiments.
@@ Appendix F Baseline Details
RoG, G-Retriever, and GNN-RAG are each trained and evaluated separately on the 8:1:1 training split of each dataset (Family, FB15k-237 and Wikidata5m) using a single NVIDIA H200 GPU, as described in Section 3.3. For RoG, we use LLaMA2-Chat-7B as the LLM backbone, instruction-finetuned on each dataset for 3 epochs separately. The batch size is set to 4, the learning rate to $2\times 10^{-5}$, and a cosine learning rate scheduler with a warmup ratio of 0.03 is adopted (linhao2024rog). For G-Retriever, the GNN backbone is a Graph Transformer (4 layers, 4 attention heads per layer, hidden size 1024) with LLaMA2-7B as the LLM. Retrieval hyperparameters and optimization follow he2024g. For GNN-RAG (mavromatis2024gnn), we use the recommended ReaRev backbone and sBERT encoder; the GNN component is trained for 200 epochs with 80 epochs of warmup and a patience of 5 for early stopping. All random seeds are fixed for reproducibility. For PoG (chen2024pog), StructGPT (jiang2023structgpt), and ToG (Jiashuo2024ToG), we use GPT-3.5-turbo as the underlying LLM, following the original papers, and adopt the original prompt and generation settings from each method. The temperature of LLMs is set to 0.
@@ Appendix G Detailed Analysis of Other Rule Types
Some rules do not exhibit simple interpretable motifs, involving unusual variable binding or rare predicate combinations. Like recursive rules (check AMIE3 lajus2020fast for more details)
@@ Appendix I Results under Permissive Evaluation Metrics
We also report the performance of representative KG-RAG models under the commonly used permissive metrics adopted in prior works, such as Hits@1 and relaxed F1/Recall/Precision (linhao2024rog; chen2024pog). These metrics treat partial string matches as correct answers and may therefore overestimate model performance, for example by counting a response that contains the gold entity in a negated form as correct (e.g., treating “not Barack Obama” as correct when the gold answer is “Barack Obama”). Nevertheless, as shown in Table 7, the results remain consistent with those obtained using our stricter evaluation protocol, revealing substantial performance differences between complete and incomplete KGs.
@@
We use ChatGPT openai2024chatgpt to enhance our writing skills, abstaining from its use in research and coding endeavors.

### v4

@@ A.1 AMIE3 Candidate Rule Refinement
AMIE3 generates candidate rules by a refinement process using a classical breadth-first search Lajus et al. (2020). It begins with rules that contain only a head atom (e.g. $\Rightarrow\texttt{hasSibling}(X,Y)$) and refines them by adding atoms to the body. For example, it may generate the refined rule:
@@ Appendix B Properties of Horn Rules Mined by AMIE3
These restrictions are widely adopted in KG rule mining (Galárraga et al., 2015; Lajus et al., 2020) to guarantee that discovered rules are logically well-formed and meaningful for downstream reasoning tasks.
@@ B.1 PCA Confidence
Definition. To address this, AMIE3 introduces PCA confidence (Partial Completeness Assumption confidence) (Galárraga et al., 2015), an optimistic variant that partially compensates for KG incompleteness. Given a rule of the form
@@
For further details, see (Galárraga et al., 2015).
@@ B.2 Rule Mining Procedure
AMIE3 generates candidate rules by a refinement process using a classical breadth-first search Lajus et al. (2020). Algorithm 2 summarizes the rule mining process of AMIE3. The algorithm starts with an initial set of rules that contain only a head atom (i.e. $\top\Rightarrow r(X,Y)$, where $\top$ denotes an empty body) and maintains a queue of rule candidates (Line 1). At each step, AMIE3 dequeues a rule $R$ from the queue and evaluates whether it satisfies three criteria (Line 5):
@@ C.1 Recall and Precision
C.2 Evaluation of Generate-on-Graph (GoG) (Xu et al., 2024)
This section provides a detailed analysis of GoG (Xu et al., 2024), following the experimental results presented in the main text. While GoG explicitly addresses knowledge incompleteness as part of its motivation, we did not include this aspect in the main body for several reasons.
GoG’s implementation is heavily optimized for specific datasets, such as CWQ (Talmor and Berant, 2018b) and WebQSP (Yih et al., 2016), and relies on several domain-specific heuristics:
@@ Appendix D Prompt Template
To ensure reproducibility and mitigate randomness in LLM outputs Potyka et al. (2024), we set the generation temperature to 0 in all experiments.
@@ Appendix F Baseline Details
RoG, G-Retriever, and GNN-RAG are each trained and evaluated separately on the 8:1:1 training split of each dataset (Family, FB15k-237 and Wikidata5m) using a single NVIDIA H200 GPU, as described in Section 3.3. For RoG, we use LLaMA2-Chat-7B as the LLM backbone, instruction-finetuned on each dataset for 3 epochs separately. The batch size is set to 4, the learning rate to $2\times 10^{-5}$, and a cosine learning rate scheduler with a warmup ratio of 0.03 is adopted (Luo et al., 2024). For G-Retriever, the GNN backbone is a Graph Transformer (4 layers, 4 attention heads per layer, hidden size 1024) with LLaMA2-7B as the LLM. Retrieval hyperparameters and optimization follow He et al. (2024). For GNN-RAG (Mavromatis and Karypis, 2024), we use the recommended ReaRev backbone and sBERT encoder; the GNN component is trained for 200 epochs with 80 epochs of warmup and a patience of 5 for early stopping. All random seeds are fixed for reproducibility. For PoG (Chen et al., 2024), StructGPT (Jiang et al., 2023), and ToG (Sun et al., 2024), we use GPT-3.5-turbo as the underlying LLM, following the original papers, and adopt the original prompt and generation settings from each method. The temperature of LLMs is set to 0.
@@ Appendix G Detailed Analysis of Other Rule Types
Some rules do not exhibit simple interpretable motifs, involving unusual variable binding or rare predicate combinations. Like recursive rules (check AMIE3 Lajus et al. (2020) for more details)
@@ Appendix I Results under Permissive Evaluation Metrics
We also report the performance of representative KG-RAG models under the commonly used permissive metrics adopted in prior works, such as Hits@1 and relaxed F1/Recall/Precision (Luo et al., 2024; Chen et al., 2024). These metrics treat partial string matches as correct answers and may therefore overestimate model performance, for example by counting a response that contains the gold entity in a negated form as correct (e.g., treating “not Barack Obama” as correct when the gold answer is “Barack Obama”). Nevertheless, as shown in Table 7, the results remain consistent with those obtained using our stricter evaluation protocol, revealing substantial performance differences between complete and incomplete KGs.
@@
We use ChatGPT OpenAI (2024) to enhance our writing skills, abstaining from its use in research and coding endeavors.